# thanks: These authors contributed equallythanks: These authors contributed equally

### v2

@@ V Discussion and Conclusions
In summary, we have provided an exactly solvable model for a non-crystalline topological order that displays qualitatively different phenomenology to its crystalline counterpart. To do so, we have extended the coupled-wire construction of the fractional Hall effect to non-crystalline systems, considering the properties of inhomogeneous arrays of coupled wires. Conservation of momentum requires an inhomogeneous system to also have inhomogeneous couplings, which ensure that different parts of the system have differently fractionalised excitations. Adjacent regions with different fractionalisation may then be coupled to one another in such a way that the system as a whole remains gapped and solvable. This is possible provided the ratio of each fractionalisation is a square of a rational number, which we enforce by setting each region’s fractionalisation to an integer $ua_{j}^{2}$. This has allowed us to propose a general and solvable construction for composite, and disordered, fractionalised systems.
@@
The phenomenology of the resulting fractionalized phase is determined by the non-uniformity, and differs significantly from the uniform case. The ground state degeneracy depends non-trivially on the fractionalisation, and consequently the wire spacing, of the constituent parts. It is given by $N_{\textit{GS}}=u\prod_{j}a_{j}$, whereas the degeneracy of a single uniform phase would be $ua^{2}$. The quasiparticles that emerge in this system also display different statistics compared to the crystalline case. We see a separation of the quasiparticles into two types: a set of lineons, which are confined to move only in the $x$ direction, and a set of composite anyons—bundles of a set number of lineons—which are free to move throughout the system in $x$ and $y$. Finally, we have calculated the anyons’ mutual statistics, finding that the composite anyons have statistics given by the common denominator of all fractions. Single lineons (which cannot braid since they are confined to a line) have mutual statistics with the composite anyons of $ua_{j}a_{j+1}$, for a lineon on the $(j+\frac{1}{2})^{\textup{th}}$ quasiwire. Finally, s-lineons live adjacent to a single wire, and braid with the statistics of quasiparticles on that wire only. Hence, both the ground state degeneracy and the excitation’s statistics offer richer possibilities compared to the uniform case.
Our work shows that correlated disorder in the wire position can lead to novel fractional quantum Hall phenomenology, in sharp contrast to the known instabilities of fractional quantum Hall edges subjected to random disorder [7, 8, 9, 10, 6, 11]. While we have not attempted to describe a microscopic Hamiltonian, it is possible to envision different experimental platforms where our construction can be relevant. It is tantalizing to speculate that strain, ever present in real-experiments, can modify the existing wire description of twisted moiré heterostructures [155, 156, 157, 158, 159, 160, 161, 162, 163] into inhomogeneous wire couplings. These could be solved using the approach presented in this work. A different approach to realize unevenly spaced wires is to use electrostatic potentials. Placing electrodes of different sizes with electrostatic potentials of alternating sign can create effective 1D wires that can couple to each other and realise our construction [175, 133]. Alternatively, heterostructures can also realize a different fractional quantum Hall effect in neighbouring layers [176]. Experiments have also demonstrated coupling between integer and fractional quantum Hall states [177, 178, 179, 180, 181] where our construction is also applicable. Finally, our model can be realised in ultra-cold atomic gases which couple 1D systems, following similar protocols as for proposals to realise crystalline coupled-wire constructions [182, 183, 184, 185, 132, 146, 186, 187].
@@
VI Acknowledgements
We thank Jens H. Bardarson, Joan Bernabeu, Alberto Cortijo, Gwendal Fève, Michele Filippone, Serge Florens, Tobias Meng, Cécile Repellin, and Pok Man Tam for helpful discussions. AGG and PD acknowledge financial support from the European Research Council (ERC) Consolidator grant under grant agreement No. 101042707 (TOPOMORPH). JS is supported by the program QuanTEdu-France No. ANR-22-CMAS-0001 France 2030.
@@ A.2 Commutation Relations
    Bare fermions bosonic field commutation relations:
@@ Appendix B Charge conservation
As discussed in Section III.2.1 the total charge of the system is not conserved when coupling two regions characterized by different fractional fillings [164, 165]. In this Appendix, we show that coupling the system to a substrate allows us to restore charge conservation while keeping the system gapped.
@@
Comparing this expression with Eq. 23, we can extract the values of $n_{1},n_{2},m_{1},m_{2}$, see Eqs. 7 and 8. According to the charge conservation condition Eq. 10, the net charge created or destroyed on each wire is $n_{j}$, and the total charge is conserved only if $n_{1}+n_{2}=0$. This condition is violated here, as $(b-a)$ charges are effectively added to or removed from the system, indicating a breakdown of charge conservation.
This issue can be resolved by introducing an additional substrate wire at the interface, serving as a charge reservoir. An example configuration is shown in Fig. S1. For fermions, $a$ and $b$ are odd, and hence $a-b$ is even, allowing for the symmetric placement of $(a-b)/2$ charges on right- and left-moving modes of the extra wire. Labelling the substrate wire with the subscript $\mathrm{sub}$, we assign $s_{\mathrm{sub}}^{R}=(a-b)/2$ and $s_{\mathrm{sub}}^{L}=(a-b)/2$, resulting in: $n_{\mathrm{sub}}=(a-b),m_{\mathrm{sub}}=0$.
@@
[…]
Importantly, this expression only depends on the bosonic phase field $\varphi_{\mathrm{sub}}$, so the substrate wire remains non-fractionalised. The coupling always adds or removes an even number of fermions to the substrate, since $a-b$ is always even. Thus, the substrate may interpreted to be in a superconducting phase where charge is conserved modulo $a-b$ [102, 188, 189].

### v4

@@ VII Discussion and Conclusions
To do so, we have extended the coupled-wire construction of the fractional Hall effect by considering the properties of spatially inhomogeneous arrays of coupled wires. Conservation of momentum requires an inhomogeneous system to also have inhomogeneous couplings, which ensure that different parts of the system have differently fractionalised excitations. Each wire is parametrised by an integer $ua_{j}^{2}$, where $u$ defines the global fractionalisation, and each $a_{j}$ determines the fractionalisation around the $j$th wire. By choosing non-uniform $a_{j}$, we are able to construct an inhomogeneous fractional Hall phase. The array of wires is coupled to form an array of quasiwires, each of which sits between two wires. At each quasiwire, the couplings can be chosen such that low-energy quasiparticles have a distinct fractionalisation, with a large freedom to choose this fractionalisation. In practice, a quasiwire between the $j$ and $j+1$ will have $ua_{j}^{2}a_{j+1}^{2}$-fractionalised excitations. Despite this, the system as a whole remains gapped and solvable. We derive the exact arrangements of the wires in real space that allows for the construction to remain well-posed. This has allowed us to propose a general and solvable construction for composite and disordered, fractionalised systems.
The phenomenology of the resulting fractionalised phase is determined by the real-space arrangement of wires, and differs significantly from the uniform case. In particular, quasiparticles in the system display a strong fractonic constraint, whereby there are completely unable to transfer between quasiwires unless grouped together into a composite excitation, motivating the description of this phase as a fractonic fractional Hall phase.
We find three classes of fracton excitaitions, single excitations are completely immobile, behaving as lineons. These can be grouped together to form s-lineons, which can only jump to one adjacent quasiwire and back. S-lineons can then be grouped together to form C-anyons, which are free to traverse the whole system. This restricted mobility can be understood intuitively as a consequence of different regions having different fractionalisation, so moving a single quasiparticle is forbidden as it does not correspond to a transfer of an integer amount of electron charge. The only mobile particles correspond to forming bundles of quasiparticles that have a total charge commensurate with the fractionalisation of two or more adjacent quasiwires. Futhermore, we show that this constraint is completely robust—that is, no perturbation can be introduced which allows for quasiparticle transport that breaks the fracton rules.
Further supporting the fractonic nature of the phase, we find that the ground state degeneracy depends exponentially on the number of wires in the system. An intriguing question for future work is whether this degeneracy requires the protection of a symmetry and how natural such a symmetry might be, or whether the degeneracy is robust.
@@
Finally, we have calculated the anyons’ mutual statistics, finding that the C-anyons have statistics given by the common denominator of all fractions, $u$. Single lineons (which cannot themselves braid since they are confined to a line) have mutual statistics with the composite anyons of $ua_{j}a_{j+1}$, for a lineon on the $(j+\frac{1}{2})^{\textup{th}}$ quasiwire. Finally, s-lineons live adjacent to a single wire, and braid with the statistics of quasiparticles on that wire only, $ua_{j}^{2}$. Hence, both the ground state degeneracy and the excitation’s properties offer richer possibilities compared to the uniform case.
@@
VIII Acknowledgements
We thank Jens H. Bardarson, Joan Bernabeu, Alberto Cortijo, Gwendal Fève, Michele Filippone, Serge Florens, Tobias Meng, Cécile Repellin, Ady Stern, and Pok Man Tam for helpful discussions. AGG and PD acknowledge financial support from the European Research Council (ERC) Consolidator grant under grant agreement No. 101042707 (TOPOMORPH). JS is supported by the program QuanTEdu-France No. ANR-22-CMAS-0001 France 2030. C.S. is supported by Vedika Khemani through the Office of Naval Research Young Investigator Program (ONR YIP) under Award Number N00014-24-1-2098.
@@ A.2 Commutation Relations
    Bare bosonic field commutation
    $[\partial_{x}\phi^{L/R}_{j}(x),\phi^{L/R}_{k}(x^{\prime})]=\mp 2\pi i\delta_{jk}\delta(x-x^{\prime}).$ (106)
  2. 2.
    Bare sum and difference bosonic field commutation relations:
@@ Appendix B Charge conservation
As discussed in Section˜III.2.1, the total charge of the system is not conserved when coupling two regions characterized by different fractional fillings [167, 168]. In this Appendix, we show that coupling the system to a substrate allows us to restore charge conservation while keeping the system gapped.
@@
Comparing this expression with Eq.˜23, we can extract the values of $n_{1},n_{2},m_{1},m_{2}$, see Eqs.˜7 and 8. According to the charge conservation condition Eq.˜10, the net charge created or destroyed on each wire is $n_{j}$, and the total charge is conserved only if $n_{1}+n_{2}=0$. This condition is violated here, as (ba)(b-a) charges are effectively added to or removed from the system, indicating a breakdown of charge conservation.
This issue can be resolved by introducing an additional substrate wire at the interface, serving as a charge reservoir. An example configuration is shown in Fig.˜S1. For fermions, $a$ and $b$ are odd, and hence aba-b is even, allowing for the symmetric placement of (ab)/2(a-b)/2 charges on right- and left-moving modes of the extra wire. Labelling the substrate wire with the subscript $\mathrm{sub}$, we assign $s_{\mathrm{sub}}^{R}=(a-b)/2$ and $s_{\mathrm{sub}}^{L}=(a-b)/2$, resulting in: $n_{\mathrm{sub}}=(a-b),m_{\mathrm{sub}}=0$.
@@
$\begin{aligned} \mathcal{O}(x)=\cos\big{(}b\varphi_{1}&-a\varphi_{2}+(a-b)\varphi_{\mathrm{sub}}\\ &+ua^{2}b\theta_{1}+uab^{2}\theta_{2}\big{)}.\end{aligned}$ (117)
Importantly, this expression only depends on the bosonic phase field $\varphi_{\mathrm{sub}}$, so the substrate wire remains non-fractionalised. The coupling always adds or removes an even number of fermions to the substrate, since aba-b is always even. Thus, the substrate may interpreted to be in a superconducting phase where charge is conserved modulo aba-b [103, 205, 206].